Plato’s analysis of falsity

A landmark in the history of logical analysis

Job van Eck

March 29, 1999

Abstract

Plato’s theory of falsity and its preliminaries, as presented in Sophistes 254d-263d, has evoked many grave criticisms: it is said to be fundamentally flawed in several respects. Yet it appears that the main origin of this view is an incorrect reading of the section on negation, which precedes the analysis of falsity. This section is interpreted as treating negative predication; in fact it treats higher order (non-)identity propositions (F is [not] G). And it is on the basis of these (non)identity propositions that the falsity of atomic first order sentences is explained. The resulting analysis turns out to be impeccable and fully adequate to the problems at issue.

Contents

1 Introduction 2

2 The problems introduced: Sophist 237-241 4 2.1 The problem of ‘not being’ . . . . . . . . . . . . . . . . . . . . . 4 2.2 The problem of falsity . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 First comment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 Solutions: Sophist 254-264 5 3.1 Analysis of the concept of not being . . . . . . . . . . . . . . . . 5 3.2 Analysis of falsity . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.3 Comment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.4 Criticisms and refutations . . . . . . . . . . . . . . . . . . . . . . 8

4 Conclusion 10

1 Introduction

Among 20th century logicians and analytic philosophers we often find an ambiguous attitude towards Plato’s philosophy. On the one hand their demands of clarity go naturally together with a tendency to soberness in ontological claims, which some deem superfluous noise or, even worse, nonsense. And the name of Plato is directly associated with such claims. For example, take his analysis of atomic sentences of the form ‘x is F’, in which a predicate is ascribed to a subject. Standardly they are read ontologically as ‘x has F’, in the sense of ‘F is a property x has’, implying ‘F is something x partakes of.’^{1} Most analytic philosophers do not accept this as an implication of ‘x is F’. Usually they take a more nominalistic stand and criticise such an analysis as being based upon the mistaken idea that the meaning of a general predicate expression lies in its referring to a property (form, characteristic). Thus they reject the very ontological implications Plato is notorious for and which constitute the fundament of important parts of his philosophy.

At the same time Plato is recognized as a master in conceptual analysis, and accordingly he is often called one of the first analytic philosophers. And rightly so. Inspired by Socrates, who, in search of ethical principles, had tried to arrive at definitions of important moral concepts, Plato took over his method of analysis, and applied it to the whole domain of philosophy. And we do not need to wonder why many distinguished analytic philosophers have paid attention to his works,^{2} especially the later dialogues, which deal with technical philosophical problems in the fields of epistemology, methodology, the philosophy of language and logical analysis: there is a real congeniality, both in the kinds of problems stated, and the way of approaching them.

However some texts are still poorly understood. A point in case is the middle part of the Sophist (237-264) where we find Plato’s treatment of the problem of falsity. The dramatic aim of the dialogue is to characterise the sophist and capture him in a definition. Theaetetus, and a guest from Elea, the main interlocutors, want to describe him as an illusionist who creates false beliefs. But the sophist is no easy quarry: he will point out that the notion of falsity is a problematic one. Someone who has a false belief or utters a false statement believes or says something that is not the case. Now, what is not the case does not exist. But how can a belief or statement represent something that is not the case, i.e. something that is not there to be represented? Can a belief or statement be false at all? This problem was in the air in the second half of the 5th century B.C. and formed a part of the climate of scepticism and relativism characteristic of the sophistic movement of the time. The view that no one can utter a false statement and that it is impossible to contradict is ascribed to Protagoras (490-420), its main representative, and it does fit in well with the subjectivism he avowed: there is no objective reality, there is only the world of subjective experience, about which no assertions of objective validity can be made. Whatever presents itself to someone, every perception and every view has an equal individual, relative validity. The opinions a person holds are true for him. “Man is the measure of all things.” The problem of falsity and the idea that one cannot distinguish false statements from true ones was one of the roots of a general distrust in the possibility of scientific, i.e. objective knowledge; and it added fuel to skepticism regarding reason and argument as a means to arrive at assertions that could claim a status of objective validity. Hence it was important not only from a logical point of view, to solve the problem.

1However, not every atomic natural language sentence can be analysed in this way. For instance, ‘John is (a) barbaros’ will be read as ‘John is not (a) Greek’, (cf. Politicus 262d-e)

which gets a totally different analysis (see further on).

2Among others H.N. Castañeda, R.M. Hare, J. Hintikka, G. Ryle, D. Wiggins.

In Sophist 237-241 Plato brings up the problems of the notions of not being and falsity and in 254-263 he works out his solutions. The text offers great difficulties, not only for a 20th century reader who is not used to the philosophical idiom of Greek colleagues from the 4th century B.C. (it can hardly be understood by a non-specialist who reads it without a commentary). Plato scholars too are faced with serious problems of interpretation. And when, as a result, they do not succeed in offering an interpretation that represents an adequate or even only coherent explanation of falsity, Plato is blamed for it, as a matter of course. Especially the passage of 254d-263d has been the butt of multifarious and grave criticisms. To mention only a few examples, Plato’s theory of negation is said to be inadequate to sustain his account of falsity;^{3} this account allegedly is vague and indefinite.^{4} It is said not to offer an adequate solution to the problem of falsity at issue, i.e. how one can say what is not the case; charity allegedly demands us to view it as a response to another (be it much less interesting) question.^{5} Plato is claimed to be subject to various confusions which are purported to yield a totally misleading account of ‘not’.^{6} However it can be shown that all criticisms up to now regarding that passage are beside the mark.^{7} The analysis of falsity we find in the Sophist and the account of negation to be derived from it are impeccable and fully adequate to the problems at issue.

In the following I will present a paraphrase of the main points of the text, give comments and deal with some of the criticisms mentioned.^{8} In order to be able to follow the train of thought we will have to stick rather closely to Greek idiom. The result will be a distorted English for which an apology is due here.

3D. Wiggins, ‘Sentence Meaning, Negation, and Plato’s Problem of Non-Being’. 268-303 in G. Vlastos ed. Plato I: Metaphysics and Epistemology. New York. 1971.

4D. Keyt, ‘Plato on Falsity: Sophist 263B’. 285-305 in E.N. Lee, A.P.D. Mourelatos, R.

Rorty eds. Exegesis and Argument: Studies in Greek Philosophy Presented to Gregory Vlastos.

Phronesis, Supplementary Volume 1, 1973

5J. McDowell, ‘Falsehood and Not-being in Plato’s Sophist’. 115-34 in M. Schofield & M.

Nussbaum eds. Language and Logos. Cambridge University Press. 1982.

6D. Bostock, ‘Plato on “Is not” ’. Oxford Studies in Ancient Philosophy 2, 1984, 89-119.

7See J.A. van Eck, ‘Falsity without negative predication: On Sophistes 255e-263d.’ Phronesis 1995. Vol. XL, 20-47. Here one may also find the justification of the way in which section 257b-263d is described in the following.

8Some historians of logic do not even take the problem seriously. M. Kneale, for instance, writes as if it is a peculiarity of some ancient Greek philosophers to recognize a difficulty here. “From the time of Parmenides with his warning against entertaining the supposition that not-being is, the Greeks had found something mysterious in negation and consequently in falsehood.” (Hereafter she only reports in one sentence that Plato solved the puzzles. See W.& M. Kneale, The Development of Logic, Oxford, 1962, p. 21.) However, the history of analytic philosophy shows that the solution to the questions at issue is anything but a matter

2 The problems introduced: Sophist 237-241

2.1 The problem of ‘not being’

The Stranger from Elea introduces the problem with the question “Do we dare to utter the expression ‘that which is totally not’?”(237b7-8) and he asks what the phrase ‘what is not’ (to me on, litterally: the not-being, that which is not) should be applied to. Apparently it cannot be applied to something, because the word ‘something’ is always used to refer to an on, something that is. So one who uses the expression talks about nothing; in fact he is not speaking at all. (i) The next step is in the material mode and runs as follows. If we want to say something about that which is not, we will not succeed: we are not allowed to ascribe an attribute to it (in Greek idiom: to attach ‘a being to a not being’, 238a7), not even a number, because it has no attribute, nor is it one or many of anything. But the only manner to introduce it or think of it would be to do so in the singular or the plural. We cannot talk about what is not without the qualifications inherent in the use of the expression. Hence the conclusion in 238c9-10: “no, it is unthinkable, unsayable, unutterable, and unformulatable in speech.”(ii) The direct implication of this, which now follows, is even worse. What we asserted in the conclusion of ii about that which is not, is contradictory. We have attached several attributes to it (unthinkable, unsayable, etc.) and that is in contradiction with what we said just now, that it has no attributes. (iii) A downright impasse. Someone who wants to inquire into what is not, cannot introduce his subject in a consistent way (as that which is not) or make statements about it and say what it is. In short, it is no subject of discourse at all.

2.2 The problem of falsity

The above passage forms the introduction to the problem of falsity, which is worked out in the following paragraphs. The trouble is that this notion of falsity is defined in terms of precisely ‘what is not’, or, in our idiom ‘what is not the case’. Theaetetus and the Stranger want to characterise the sofist as an illusionist, someone who creates false beliefs. Now, “a false belief will be a matter of believing things that are contrary to the things that are the case”, or “believing things that are not the case.” In the same way, to make a false satement is to say things that are not the case. But, the sophist will argue, this definition implies that there are no false statements. A false statement says what is not the case. But there are no things that are not the case: a non-fact, so to speak, is not something. So a false statement is not about something. But then, it is no statement at all. If Theaetetus and the Stranger claim that there are false statements, they will be forced to do something that (as they agreed in 238a) was forbidden: to attach “being to not being”, i.e presuppose that there are facts that are not the case. Because it is a statement it is about something, a “being”, something that is the case, a fact; and because it is false, the very fact it is about is not the case. This problem holds for both types of falsity distinguished at 240e10-241a1: “And a false statement is to be regarded ... as stating of things that are, that they are not (a), and of things that are not, that they are (b)”. To give an example that will be used in 263, suppose Theaetetus is sitting. Then ‘Theaetetus is flying’ is a statement of type b, and ‘Theaetetus is not sitting’ is one of type a.

of course. Let me give some examples. In “Some Main Problems in Philosophy”, Londen 1953, G.E. Moore struggled with the problem of false opinion. B. Russell found himself forced to accept the existence of negative facts (see Monist 28, 1918 Lecture III). In “Are there propositions?” Proceedings of the Aristotelian Society, XXX, 1929-1930, G. Ryle denied that false sentences have meaning. And as late as 1980 Gochet could write about “the enigma of the meaning of false sentences” (P. Gochet, Outline of a Nominalist Theory of Propositions.

Dordrecht, 1980, ch. VIII). (The references to Moore and Ryle are borrowed from this work.)

2.3 First comment

What immediatelly will strike the reader in this argument is the transition from ‘saying something’ to ‘being about something’. In the Greek original this is concealed because for both notions the same phrase is used: legein ti, which may mean both saying something - in the sense of asserting something - and talking about, discussing something. Another point is the use of the verb einai, ‘being’. If it does not occur with any further adjunct it may be elliptical (a complement has to be understood, for instance in the form of an adjective, or a noun, or a participle), but it may also be complete in itself, having the meaning of being the case, being real, existing. Thus the problem of falsity is easily formulated in Greek idiom. However, the argument on which it is built is not dependent on this idiom. At its root lies a notion of language, that, irrespective of any idiomatic particularity, at first sight is so natural that it may more or less inadvertently be adhered to as a matter of course, viz. the idea that the meaning of a linguistic expression (a word, term, phrase, sentence) is that which it applies to. Thus a sentence would be a kind of complex name of what it is about, the state of affairs it bears upon. (How self-evident this idea seems to be also appears from the fact that an acute logician like B. Russell has been a victim of it. “... I had never realised it [that propositions are not names of facts] until it was pointed out to me by a former pupil of mine, Wittgenstein.”)^{9}

3 Solutions: Sophist 254-264

3.1 Analysis of the concept of not being

In 242 the man from Elea directs his attention to the notion of being, which causes problems as well. Earlier philosophers had not been clear about it, and both the materialists and the ‘friends of the forms’ have one-sided views, the last mentioned because they are of the opinion that only that which is always at rest and never suffers change is something. It is, however, not only rest, but also change that has to be recognized as a way of being. In 254d-255a it is pointed out that being, rest and change do make three forms, which implies that each is different from the other ones; thus difference emerges and is introduced now, together with sameness as a possible fourth and fifth form. Then it is argued that sameness and difference are two other forms indeed, distinct from the first three. In 255e the Stranger takes the form change and examines what its relation to the other forms is. He shows that we can call change the same and not the same without contradiction. In calling it the same we say that it has the attibute the same, it partakes of sameness (i.e. with respect to itself, it is identical to itself); and at the same time we can say that it is not the same, in that change is different from (the form) the same (change ≠ sameness). This is generalized: we have a distinction between an ‘is’ as in ‘is F’, due to participation of F, and an ‘is’ as in ‘is not F’, due to participation of difference from F. We can even say that change is being and not being: it partakes of being and thus is something, and it is different from being. Thus it turns out that not being (X), in the sense of difference from being (X) really is something. And everything participates of it; even being itself is not (X) insofar as it is different from X.

9B. Russell, The Philosophy of Logical Atomism. Monist 1918-19, Lecture I.

This is a breakthrough of the impasse resulting from the argument of 237-239 that no being could be attributed to not being. Mind you, the argument has not been refuted. The point is that the notion of not being it is about, is that of totally not being, being nothing at all, in other words, the opposite of being. What has been shown now is that something can be not X without being the opposite of X, and so can be not being without being the opposite of being: when we say that something is not X we put it in contrast to, set it against X. This setting against X is only difference from X, not the opposite of X. The argument of 237-239, involving the notion of totally not being, being nothing at all, still holds, but it does not apply to the new notion of not being. “We have said goodbye long ago to any opposite of what is, and to whether it is something or not, and also to whether or not an account can be given of it.” (258e8-259a1).

So the definition of falsity must be understood in the light of the analysis of not being just arrived at. Has the problem of falsity been resolved now? No. The sophist has another barrier to put up. The question whether somebody can say ‘things that are not the case’ (me onta, not-beings) has now become whether what he says can be different from the onta, from what is the case. The sophist challenges this (260d-e); the reason why is easy to reconstruct. Every statement is about something, is the representation of something, a state of affairs. It had already been observed (237e1-5) that this cannot be totally nothing: if the statement were about nothing the speaker would not say anything at all. If there is something wrong with the statement it can be at most that it is a representation of another case, some different thing (cf. Euthydemus 286a4-b6 on the impossibility of contradiction). But, the sophist will continue, this is not the same as its being about something different from what is the case. The other thing is also something and so cannot be something else than what is the case. Therefore, that which poses as a false statement about something is either a true statement about something else, or no statement at all. Saying something different from what is the case is not possible, because what is different from what is the case does not exist, is utterly not being.

3.2 Analysis of falsity

In 261c6-262e1 the interlocutors bring up the question what the structure of a logos (sentence, statement) is. First the Eleatic stranger (ES) distinguishes between two kinds of words, each designating something, i) rhemata (verbs, attributes) indicating states or actions, ii) onomata (names, including nouns) designating the subjects of these states and actions. A statement never consists solely of a string of names, such as ‘lion, stag, horse’, nor of a series of verbs without names, like ‘walks, runs, sleeps’. This does not yet make a statement: to get a statement you have to combine a verb with a name. “The moment you do that they fit together and the first combination straight off becomes a statement, of the simplest and smallest kind you might say.” (262c5-7) Someone who says ‘Man understands’ does not merely name something, he states something, and in fact it is this complex of words we use the word ‘statement’ for.

Now (263) the ES is ready for an analysis of a false logos. He takes two sentences, ‘Theaetetus sits’ and ‘Theaetetus flies’ and remarks that both are about something: Theaetetus. One is true, the other false. The discussion continues: “ES: (263b4-5) And the true one says of the things that are that they are about you. Th: Certainly. ES: (263b7) Whereas the false one says so of other things than the things that are. Th: Yes. ES: (263b9) And accordingly it speaks about things that are not as things that are.” Then, recapitulating, the ES points out again that the false sentence does have a subject: Theaetetus. If it were not ‘of something’, i.e. did not have something as a subject it would not be a statement at all. And he concludes that an elementary false statement is a make-up of names and attributes (verbs) in which things are stated about a subject, but in such a way that “what is different is stated as what is the same, and what is not the case as what is.” (263d1-2)

3.3 Comment

The best known part of this is found in 262, the distinction between the naming level, and the stating, or predicating level: both the name and the attribute designate (name) something (a thing and a state or action, respectively), but only the logos as a whole, consisting of a name plus an attribute, says, states something by saying that what is indicated by means of the attribute is the case about that which is referred to by the name.^{10} And when the ES wants to say that the statement ‘Theaetetus flies’ is about something he does no longer use the ambiguous ‘a statement legei ti’- says (?), discusses(?) something- he says that a statement is ‘of something’.

10See G. Nuchelmans Theories of the Proposition. Amsterdam, 1973, 14-18)

This opens the possibility to solve the last problem the sophist had raised: how can a statement say something that is different from what is the case? Let us go back to 263b4-5. Litterally the text reads “And the true one says of the beings that they are about you” So sitting is a being about Theaetetus and of that sitting “Theaetetus sits” says indeed that it is (the case) about Theaetetus. Instead of ‘being the case’ we have now ‘being the case about ...’ (compare the now current semantic expression ‘is true of ...’ , as used in the phrase ‘Predicate F is true of object x’). 263b7: “Whereas the false one says so of other things than the things that are”. So “Theaetetus flies” says that flying is the case about Theaetetus, whereas flying is something else than (different from) the things that are in fact about Theaetetus. (With 263b9 the ES points out that “Theaetetus flies” answers to the description of a false sentence we met in 241a1).

In this way a logos can say something that is different from what is the case. That which in a logos is said to be the case, has been analysed as something that is said to be the case concerning the subject; thereby the monolithic character of its content, the on it is supposed to be about (in our examples, Theaetetusbeing-in-flight, Theaetetus-being-sitting, respectively), is taken away. The aspect of difference accounting for the falsity of the sentence does not bear upon the whole, undivided state of affairs ‘Theaetetus-being-in-flight’, which is a total not being indeed, only upon what is said about Theaetetus: being in flight. The ‘other things’ are not in contrast to facts, to onta straight, but in contrast to onta which may be about something, in short, to being this or that.

3.4 Criticisms and refutations

Further comments may best be given against the background of some of the criticisms mentioned earlier. For instance, it is said that this analysis of falsity does not show how to cope with the problem of falsity as stated above. The point at issue in the problem of the falsity of ‘Theaetetus is in flight’ is not the thing the ES is talking about, the attribute ‘in flight’ that is not in relation to Theaetetus. The problem concerns the unavailability of the flight Theaetetus is in; and that is a different item. The falsity of ‘Theaetetus is in flight’ consists in the fact that the state of affairs the sentence represents and the aspect that corresponds to the rhema, that is, Theaetetus’ flight, do not exist and are utterly not, in the sense of the contrary of being. In consequence, the arguments in 2376b-239c8 which disparage this notion still undermine the possibility of falsity.^{11}

In fact it is easy to reply on behalf of the ES. Sure enough, that Theaetetus is flying is not the case, it is so to speak a nonfact, and Theaetetus’ flying does not exist. But the sentence “Theaetetus flies” is not about that, not about a nonexistent state of affairs, a “nonfact”, and not about Theaetetus’ flying. And then, in fact, that is not what makes it a false sentence. What the ES did show is that there is another way in which you can describe what a falsity consists in, what it is that you say when saying what is not the case. That “Theaetetus flies” states “the things that are not as being” (263b9) does not imply that this sentence speaks about a fancied, unreal fact as a fact, or about Theaetetus’ flying as if it existed (were the case), but that it is about flying in general, which, in contradistinction to the “utterly not being” flying of Theaetetus, is something indeed, only different from what Theaetetus in fact is doing. In a word, it speaks about something that is not the case concerning Theaetetus (in that it is different from everything that is concerning him) as something that is the case concerning him. Thus, a statement can be false without being about a nonexistent situation or an individual action not done. If one would like to say at all that in false statements the subject of discourse is what is not the case, or that nonexistent states of affairs are represented in them, then the ES does not allow us to take this in the strict sense of: false statements are about that which is not the case, but in the sense of: a false statement makes it appear as if a nonexistent state of affairs does exist, without literally taking as a subject that which does not exist, but by speaking, in the way mentioned before, about a relative not being (which does exist) as being.

11See J. McDowell o.c.

Another point is that some commentators think that the ES needed or presupposed for his analysis of the falsity of ‘Theaetetus is flying’ the notion of ‘not being’ as involved in statements of negative predication like ‘Theaetetus is not flying’. Then they perceive a treatment of negative predication in 257c-258c, a fatal mistake nearly all commentators make and one of the main sources of the criticisms mentioned earlier: if one interprets what is said there in terms of negative predication one can hardly make sense of the passage. For instance, according to David Wiggins (o.c.) the ES suggests in 257b ff. that in statements of the form ‘x is not F’ a negative property is attributed to x, introduced by the expression of the form ‘not F’, which is taken as one phrase. He shows the disastrous consequences of such an analysis and concludes that very probably there is a grave confusion on Plato’s part. Should we impute this view to Plato? It must be admitted that Plato does take a phrase of the form ‘not F’ as one ‘rhema’(257b7), that introduces a negative property, which he indicates as ‘some of the (things) other than F.’ In 257b9-c1 we find: “When a denial is said to signify an opposite, we shall not agree, but only on this, that the prefixed ‘not’ indicates some of the things other than ... the things the words, uttered after the denial, stand for.” Thus it seems that for Plato a statement of the form ‘x is not F’ is the assignment of some negative property to x, which commentators feel forced (in the light of 257b10) to read as ‘x has some property different from F’. A sadly inadequate analysis of negative predication. Some interpret ‘different from F’ here as ‘incompatible with F’, thus presupposing a sudden but unannounced shift in sense of ‘different’ to ‘incompatible’ (or at least a crucial restriction to a range of incompatible properties) on the part of the ES.^{12}

12This interpretation is defended e.g. by F.J. Pelletier, Parmenides, Plato, and the Semantics of Not-Being. The University of Chicago Press, 1990.

However, we do not find any treatment of negative predication in 257- 258. The sentence “ ‘not’ indicates some of the things other than ... the things the words, uttered after the denial, stand for” is not meant to determine the meaning of a negative phrase ‘not F’ - it does not say ‘not F’ = (by def.) ‘some other thing than F’. It says that for the meaning of ‘not F’ we have to look (not at something opposite to F, but only) at some other thing than F, and this other thing turns out to be difference from F. The function of the passage is to explain ‘the nature of’ not being F, and to show that it is not the opposite of F. And it does so still in the context of higher order statements such as ‘(the property, form) X is not F’, meaning: the form X is different from the form F, or: partakes of difference from F. Thus not F can indeed be viewed as a negative property, i.e. the second (or higher) property difference from F. As long as we recognize that 257b1-258e3 is not meant as an analysis of negative predication, we see that the placing of not before a predicate F, taking not F as a rhema that refers to a (negative) property, does no harm at all.

In fact, there is no trace of negative predication in the analysis of falsity in 263 either, neither is it supposed, nor needed. In ‘Theaetetus flies’ “things are stated about a subject, but in such a way that what is different is stated as what is the same, and what is not as what is” (263d1-2). Thus we see in which way, in ‘Theaetetus flies’, what is not is stated as what is. It is not that Theaetetus’ not being in flight (negative predication) is described as the flying of Theaetetus, but that things that are different from what is the case concerning him (viz. flying) are described as the same (as what is the case about him).

4 Conclusion

Now, it may be surprising to find atomic first order propositions explained by way of higher order (non-) identity propositions in this way and to see difference and sameness, as a contradictory pair, function as the primitives in terms of which falsity is defined.^{13} But it is pointless to complain, as White does,^{14} that Plato’s treatment of negation ‘cuts off the natural path’ of explaining the relation between ‘the same’ and ‘different’, leaving it primitive. In defining falsity or negation one must start somewhere, and as ‘the same’ and ‘different’ are Plato’s conscious starting point this criticism is just as preposterous as saying that the semantics of predicate logic falls short because it cannot explain the relation between inclusion in ^{15} and exclusion from a class. Indeed, from a logical point of view the couple different-the same, as a pair of primitives, is on a par with exclusion-inclusion: the nature of the relation between difference and sameness (their being contradictories) is no less clear than that between being excluded from a class and being included in it; and to say that x belongs to the class of F-things or classifying x as an F-thing boils down to saying that F is a characteristic of x, or identifying one of the characteristics of x as F. And for a Platonist it is even the more fundamental pair from an ontological point of view: as x belongs to the class of F-things because F is a characteristic of x (and not the other way around), in the statement ‘x is F’ basically a characteristic of x is identified as F. The identification Plato’s analysis of ‘x is F’ as ‘F is a form x participates in’ involves, comes out more clearly if we represent the statement in a formula. If we use the capital ‘Y’ as a first order variable pertaining to forms (properties, characteristics), we have ∃Y(Yx∧Y=F): there is something x participates of, which is (identical to) F.

13In the falsehood arising from the negation of a true atomic logos (Theaetetus is not sitting)

ta auta -things that are the same (as etc.)- are described as thatera -things that are different (from etc.). In the true negation ‘Theaetetus is not flying’ thatera are described as thatera, in the true statement ‘Theaetetus is sitting’ ta auta as ta auta. These analyses can be derived by adapting the wordings of 263d1-2 to the possibilities of falsity proposed in 240e241a. See van Eck, o.c. 43-44.

14See N.P. White, Plato. Sophist. Translated with Introduction and Notes. Cambridge, 1993.

15For the sake of convenience I use the term inclusion here as a contradictory of exclusion.

Normally it stands for a relation between classes, represented by ⊆.

If this Platonic stand is taken the solution of the problem of falsity is easy: with the above analysis it is given. ‘x is F’ is false if there is not something x participates of and which is identical to F, in other words, everything x does participate of is different from F. The only ‘negative fact’ this presupposes,^{16} that there is a property different from F, boils down to there being more than one property, and F being one of them.

16On the supposition, of course, that ‘x’ does refer to an object having (one or more)

properties.